# Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

### v2

@@ 5.2 Experimental Results
Condition Dim 1 Dim 2 Dim 3 Dim 4
w/o context 0.672-0.672 0.384-0.384 0.425-0.425 $0.584$
w/ context $0.097$ $0.728$ $0.165$ 0.174-0.174
5.3 Further Analysis
Q1: Do LLMs Employ Correct Symbolic Reasoning Paths? While previous research Hao et al. (2024) has demonstrated that yielding correct answers does not necessarily yield correct reasoning paths, we find that general LLMs like Qwen2.5-14B with pure textual input are able to perform a certain level of reasoning (through CoT prompting) over the time series and the given context. As shown in Table H.2, the LLM is not only able to consider the meaning of each variable, but also the historical candidates.
Q2: How should test-time compute be structured for consistent improvement? While scaling test-time compute can enhance LLM performance Snell et al. (2024), our findings indicate that the structure of this computation is critical for achieving consistent improvement. Simply increasing compute via naive CoT reasoning does not yield reliable gains, as shown in Obs. 3. This is because the depth of naive CoT reasoning often remains shallow and lacks key mechanisms such as reflection, verification, and backtracing that are present in more advanced RLM Besta et al. (2025) with Long CoT Chen et al. (2025), which involves more cognitive behaviors such as verification, reflection, and backtracing, etc. We use illustrative examples to demonstrate this in Appendix H.3.
(a) Comparison of increasing test-time compute on CDEs.
(b) Comparison of w and w/o thinking on BNs and SCMs.
Figure 3: Comparison of performance with and without thinking.
Method $SR_{ID}^{2}$ $SR_{OOD}^{2}$ Sym. Prox. $\downarrow$ Comp. $\downarrow$
GPLearn 27.6 14.9 6.083 1.333
GPT-4o-mini 39.3 24.1 7.438 7.406
GPLearn + LLM-as-Judge 31.7 16.3 6.550 1.200
LLM-as-Predictor + GPLearn 89.5 69.3 5.045 2.682
Method Setting Metrics
Prec. Recall F1 Acc BM Comp. $\downarrow$
LogicGep ID 88.0 88.6 88.0 91.7 80.9 0.773
OOD 78.9 83.0 80.0 85.6 69.2 0.773
GPT-4o-mini ID 52.8 65.6 57.8 61.2 23.6 2.422
OOD 49.9 62.3 54.5 58.3 17.3 2.422
LogicGep + Judge ID 89.3 92.8 90.9 93.2 85.6 1.012
OOD 78.6 85.2 80.4 85.8 71.0 1.012
Predictor + LogicGep ID 72.2 84.7 77.5 80.6 61.0 0.993
OOD 63.1 77.2 67.9 73.2 46.9 0.993
Q3: Does the generalizability of prediction correlate with structural complexity? When faced with multiple symbolic structures that can fit the given time series, the conventional wisdom suggests that the simplest one is the most likely to be generalizable. Existing LLM-based approaches for symbolic regression Shojaee et al. (2024); Li et al. (2024); Merler et al. (2024); Grayeli et al. (2024); Wang et al. (2025), rooted in the principle of Occam’s Razor, advocate for choosing the expression with the least complexity. However, our experimental results challenge the assumption that simpler is always better. Our analysis shows that expressions with higher complexity can also achieve improved symbolic proximity and OOD performance in Table 2. This is because, in complex scientific domains such as biology, physics, and healthcare Xie et al. (2025), the ground-truth expression may inherently be more structurally complex while being numerically simpler in the ID setting. Imposing a strong simplicity criterion can hinder the discovery or selection of the more generalizable context-aligned candidates. As detailed in Appendix E and Table 5, our analysis suggests that: Candidate ranking and selection should not solely rely on complexity and may benefit from considering contextual information.
Q4: Can LLMs and genetic programming (GP) be combined for enhanced performance? While LLMs show superior performance on CDE and SCM inference, GP holds value for its cost, control over complexity, and stronger performance in BN inference. To capitalize on the advantages of both, we explored two primary methods for creating a high-performance hybrid system, as detailed in Appendix F. The key is to use the contextual understanding of LLMs to “inject context and knowledge” into the GP workflow, boosting its performance. This can be achieved in two main ways: (a) LLM-Guided Initialization: Use an "LLM-as-Predictor" to generate a high-quality initial population of candidate expressions for the GP to evolve. This provides the GP with a context-aware and promising starting point. (b) LLM-Guided Evaluation: Employ an “LLM-as-Judge” to provide context-enhanced evaluation scores during the GP’s fitness assessment. This helps guide the evolutionary search toward solutions that are not only numerically accurate but also scientifically plausible. As shown in Tables 6 and 7, combining LLMs with GP in these roles demonstrably improves performance.
6 Conclusion and Outlook
Future Opportunities. Based on our observations and takeaways, we summarize some potential future directions as follows: (i) task-specific scaling of test-time compute and reasoning depth; (ii) richer, knowledge-heavy context to guide hypotheses; (iii) context-aware criteria beyond syntactic simplicity (e.g., plausibility, robustness, causal faithfulness); and (iv) broader symbolic targets plus verifiable reasoning traces for interpretability.
7 Limitations
This work has several limitations stemming from computational constraints. First, we did not evaluate larger open-source models (e.g., DeepSeek-R1). Second, we capped the budget at 100 generation epochs and a maximum of 20 retries per epoch for each run; samples without candidates that could reach the tolerance will be rerun once again to mitigate uncertainty. Finally, we examined only one LLM and one GP in the hybrid architecture, and did not analyze the effects of reasoning strategies or model size.
Appendix A Related Work
LLM Symbolic Reasoning. Due to the strong in-context learning and reasoning ability of Large language models (LLMs) that allow them to adapt to various tasks, they are able to perform logical inferenceAhn et al. (2024); Wang and Chen (2023), and temporal symbolic reasoning Fang et al. (2024). Despite the strong ability of LLMs, current research for SR tasks only applies them as proposal generators without a reasoning process. Though the generated functions may achieve a high fitting score, the reasoning process remains unknown to us and may fail to align with the context, with little real-world meaning. In addition, the current research focus has also skewed toward algebraic equations, with little work on logic rules or causal relations. Our SymbolBench, compared to previous benchmarks as shown in Table 1, addresses this by systematically evaluating both LLMs and MLLMs on SR for time series data, covering varied symbolic forms and emphasizing the transparency of the reasoning process. In addition, compared to prior benchmarks, we provide a more comprehensive evaluation including In-distribution and Out-of-distribution settings over coupled ODEs, logical expressions, and structured causal models.
Appendix B SymbolBench Dataset
dx2/dt = (c8 + c9 * x3) - c10 * x2 x2: microg_dl
dx3/dt = (c4 + c5 * x1) / (1 + c6 * x2) - c7 * x3 x3: pg_ml
x2 = ( x1 AND NOT ( ( x3 OR x5 ) OR x4 ) ) x2: v_Emx2
x3 = ( ( x3 AND x5 ) AND NOT x2 ) x3: v_Fgf8
x4 = ( x5 AND NOT ( x2 OR x1 ) ) x4: v_Pax6
x5 = ( x3 AND NOT x2 ) x5: v_Sp8
$\sim$ x1: rapidly_activating_K_current_n_gate (dimensionless)
$\sim$ x2: Ca_i in component ionic_concentrations (micromolar)
$\sim$ x3: None

### v3

@@ 5 Experiment
Avg. ID Rank $\downarrow$ $17.0$ $13.0$ $16.0$ $3.5$ $1.0$ $2.0$ $3.5$ $15.0$ $14.0$ $10.5$ $9.0$ $6.5$ $9.5$ $9.5$ $5.0$ $9.5$ $7.5$
Avg. OOD Rank $\downarrow$ $17.0$ $8.0$ $16.0$ $7.0$ $13.0$ $14.0$ $1.5$ $11.0$ $13.0$ $3.0$ $1.5$ $8.5$ $7.5$ $8.5$ $5.0$ $7.0$ $10.5$
5.1 Experimental Setup
5.2 Experimental Results
Across all three tasks, we select LLMs with various sizes and architectures, including Qwen2.5-14B Team (2024), Llama-3.2B Dubey et al. (2024), Mathstral-7B Jiang et al. (2023), GPT-4o-mini Achiam et al. (2023), and ChatTS-14B Xie et al. (2024). For GPT-4o-mini and ChatTS-14B, we use inputs with visual and temporal modalities. We present the results in Table 2,  3, and  4, where ablation studies on the four prompts with Qwen2.5-14B are integrated into the main tables rather than reported in a dedicated section.
Obs. 1: LLMs demonstrate superior capability compared to baselines on CDEs and SCMs datasets, while failing to compete against baselines in Boolean network inference. For Boolean network inference, while LLMs consistently achieve positive bookmaker informedness scores, indicating performance better than random guessing, the genetic programming-based model, LogicGep, significantly outperforms all evaluated LLMs across nearly all evaluation metrics in both ID and OOD scenarios. This could be explained by both the symbolic and numerical fitting process. (a) Symbolic fitting Unlike CDEs, BN inference does not involve coefficient optimization. Compared with CDEs, even if the symbolic structure is not precisely correct, adjusting coefficients can still yield good numerical accuracy. Compared with SCMs, SCM inference focuses only on discovering causal relationships, which is a less stringent goal than recovering the full dynamics as in BN inference. (b) Numerical fitting: CDEs and SCMs are inferred from continuous time series, which form a consistent chains of state transitions, whereas BN inference uses state transitions with various initial conditions, forming graphs of state transitions, as shown in Appendix B and  J.1, which can be hard to identify trends and summarize patterns. Further error analysis is provided in Appendix F.
Obs. 3: Chain-of-thought prompting does not consistently improve performance. Except ChatTS-14B, further introducing CoT prompting does not lead to consistent gains, especially on the CDEs dataset. While increasing test-time compute has been shown to improve outcomes in various tasks Snell et al. (2024), to which CoT can contribute, similar limitations are observed in other benchmarks like SciBench Wang et al. (2023).
Obs. 4: Providing problem contexts improves LLM performance. Across all three tasks, compared to the Naive prompt, LLMs demonstrate the ability to leverage the provided context, resulting in higher numerical performance. In this setting, context functions as a form of conditioning, helping to constrain the solution space and guide the model toward more accurate and relevant inferences with a faster convergence rate, as shown in Appendix I.
Table 3: Comparison of Boolean network inference across ID and OOD settings. Full results in Appendix E.
Model Setting ID OOD Symb. Prox. $\downarrow$ Comp. $\downarrow$
Table 4: Comparison of causal discovery methods and LLM-based approaches. Full results in Appendix  E.
[…]
PCMCI $\sim$ 52.7 46.1 23.7 7.9 95.28
LPCMCI $\sim$ 52.0 29.1 18.4 5.3 25.35
j-PCMCI+ $\sim$ 46.2 39.1 13.2 7.4 49.36
Qwen2.5-14B Naïve 49.7 40.6 17.9 8.4 35.54
Base 50.5 38.6 18.4 8.9 34.58
Context 53.4 37.0 19.5 7.9 31.69
CoT 51.3 38.1 20.0 8.4 39.36
ChatTS-14B Context 54.1 27.7 22.6 11.6 32.48
CoT 54.4 27.7 25.3 10.5 31.10
Condition Dim 1 Dim 2 Dim 3 Dim 4
w/o context 0.672-0.672 0.384-0.384 0.425-0.425 $0.584$
w/ context $0.097$ $0.728$ $0.165$ 0.174-0.174
5.3 Further Analysis
Q1: Do LLMs Employ Correct Symbolic Reasoning Paths? While previous research Hao et al. (2024) has demonstrated that yielding correct answers does not necessarily yield correct reasoning paths, we find that general LLMs like Qwen2.5-14B with pure textual input are able to perform a certain level of reasoning (through CoT prompting) over the time series and the given context. As shown in Table J.2, the LLM is not only able to consider the meaning of each variable, but also the historical candidates.
Q2: How should test-time compute be structured for consistent improvement? While scaling test-time compute can boost LLM performance Snell et al. (2024), our results show that the structure of scaling is crucial for consistent gains. Simply applying naive CoT reasoning offers limited benefit (Obs. 3) due to shallow reasoning depth and the absence of mechanisms like reflection and verification. In contrast, advanced RLM Besta et al. (2025) with Long CoT Chen et al. (2025) incorporates these cognitive processes, leading to more robust improvements. Illustrative examples are provided in Appendix J.3.
(a) Comparison of increasing test-time compute on CDEs.
(b) Comparison of w and w/o thinking on BNs and SCMs.
Figure 5: Comparison of performance on thinking.
Method $SR_{ID}^{2}$ $SR_{OOD}^{2}$ Sym. Prox. $\downarrow$ Comp. $\downarrow$
GPLearn 27.6 14.9 6.083 1.333
GPT-4o-mini 39.3 24.1 7.438 7.406
GPLearn + LLM-as-Judge 31.7 16.3 6.550 1.200
GPLearn + LLM-as-Predictor 89.5 69.3 5.045 2.682
Method Setting Metrics Comp. $\downarrow$
Prec. Recall F1 Acc BM
LogicGep ID 88.0 88.6 88.0 91.7 80.9 0.773
OOD 78.9 83.0 80.0 85.6 69.2
GPT-4o-mini ID 52.8 65.6 57.8 61.2 23.6 2.422
OOD 49.9 62.3 54.5 58.3 17.3
LogicGep + Judge ID 89.3 92.8 90.9 93.2 85.6 1.012
OOD 78.6 85.2 80.4 85.8 71.0
Predictor + LogicGep ID 72.2 84.7 77.5 80.6 61.0 0.993
OOD 63.1 77.2 67.9 73.2 46.9
Figure 6: Performance degradation on the CDE (Left, Middle) and SCM (Right) datasets after introducing Gaussian noise to the time series.
Q3: Does the generalizability of prediction correlate with structural complexity? When multiple symbolic structures fit a time series, conventional wisdom, rooted in Occam’s Razor, favors the simplest as most generalizable. Recent LLM-based symbolic regression methods Shojaee et al. (2024); Li et al. (2024); Merler et al. (2024); Grayeli et al. (2024); Wang et al. (2025) follow this principle. However, our results challenge the notion that simpler is always better. As shown in Table 2, more complex expressions can yield higher symbolic proximity and OOD performance. In domains like biology, physics, and healthcare Xie et al. (2025), true underlying laws may be structurally complex yet numerically simple in-distribution. Overemphasizing simplicity can therefore obscure more generalizable, context-aligned candidates. As detailed in Appendix G and Table 5, our findings indicate that: Candidate ranking and selection should not solely rely on complexity and may benefit from considering contextual information.
Q4: How do LLMs perform with noisy data? To evaluate robustness to real-world data imperfections, we introduce Gaussian noise at varying levels (0.01 to 0.4) to the time series data and assess performance degradation across different model scales. We sample 20 cases for each experiment. As shown in Figure 6, on the CDE dataset, most models remain vulnerable to the increasing noise, especially in the OOD setting. On the other hand, on the SCM dataset, larger models exhibit increasing robustness. Overall, all models experience performance degradation, and the impact varies across tasks.
Q5: Can LLMs and genetic programming (GP) be combined for enhanced performance? While LLMs show superior performance on CDE and SCM inference, GP holds value for its cost, control over complexity, and stronger performance in BN inference. To capitalize on the advantages of both, we explored two primary methods for creating a high-performance hybrid system, as detailed in Figure 3 and Appendix H. The key is to use the contextual understanding of LLMs to “inject context and knowledge” into the GP workflow, boosting its performance. This can be achieved in two main ways: (a) LLM-Guided Initialization: Use an "LLM-as-Predictor" to generate a high-quality initial population of candidate expressions for the GP to evolve. This provides the GP with a context-aware and promising starting point. (b) LLM-Guided Evaluation: Employ an “LLM-as-Judge” to provide context-enhanced evaluation scores during the GP’s fitness assessment. This helps guide the evolutionary search toward solutions that are not only numerically accurate but also scientifically plausible. As shown in Tables 6 and 7, combining LLMs with GP in these roles demonstrably improves performance.
6 Conclusion and Outlook
Future Opportunities. Scientific discovery from observations lies at the core of human science development. Focusing on one specific type of observation: temporal data, we highlight several opportunities to push the development of automation in scientific discovery: (i) task-specific scaling of test-time compute and reasoning depth; (ii) knowledge-heavy context and context-aware verifications beyond syntactic simplicity to guide hypotheses; (iii) adaptation to diverse multimodal observations, and (iv) integration with efficient Genetic Programming algorithms.
7 Limitations
This work has several limitations stemming from computational constraints. First, we provide evaluations on large open-source models (e.g., DeepSeek-R1) only in Figure 5 instead of in the main tables. Second, we capped the budget at 100 generation epochs and a maximum of 20 retries per epoch for each run; samples without candidates that could reach the tolerance will be rerun once again to mitigate uncertainty. Finally, we examined only one LLM and one GP in the hybrid architecture, and did not analyze the effects of reasoning strategies or model size.
Appendix A Related Work
LLM Symbolic Reasoning. LLMs exhibit strong in-context learning and reasoning capabilities, enabling them to perform tasks such as logical inference Ahn et al. (2024); Wang and Chen (2023) and even preliminary forms of temporal reasoning Fang et al. (2024). Prior work on time series reasoning with LLMs, however, has largely emphasized linguistic outputs, for example, generating natural-language summaries of temporal patterns, identifying anomalies, or extracting descriptive insights from sequential data. While valuable, these approaches focus on surface-level interpretation and do not attempt to uncover the underlying symbolic structures that govern dynamic systems. In contrast, our work targets symbolic reasoning tasks where the output is not a narrative but a formal structure: scientific equations, logical rules, or causal graphs. These forms require abstraction, generalization, and alignment with domain context, going beyond function generation. Existing efforts to apply LLMs to symbolic regression have mainly treated them as equation generators Merler et al. (2024); Li et al. (2024); Shojaee et al. (2024), producing functions that may fit data but often lack explanatory power or contextual plausibility. Moreover, the symbolic space considered so far has been restricted almost exclusively to algebraic expressions, leaving other important forms such as Boolean networks Zhang et al. (2024) and causal relations Assaad et al. (2022) largely unexplored.
Appendix B SymbolBench Dataset
To rigorously evaluate the reasoning abilities of LLMs on time series-related science discovery, we introduce SymbolBench, a curated benchmark that aims to uncover symbolic structures from time series. The dataset spans diverse domains (e.g., physics, biology) and is structured around three core categories of symbolic structures.
Coupled Differential Equations (CDEs). CDEs represent dynamic systems, yielding continuous and multivariate time series data. Let $x_{i}$ denote the $i$-th variable of a multivariate time series, a coupled differential equation can be described as the following: $\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{n}),i=1,...,n$, where the symbolic structures $f_{i}$ describe dynamic interactions among state variables. The corresponding time series data consists of numerical solutions $\{x_{i}(t)\}$ over time, generated from initial conditions and system parameters. This setting reflects real-world dynamical systems in physics and engineering, where the complexity arises from variable interdependence. While the previous benchmark dataset, ODEbench d’Ascoli et al. (2023), contains coupled ODEs from 1 dimension to 4 dimensions, the number of samples is small, and the class of dimensions is heavily imbalanced, with only three 4-dimensional ODEs. In this study, we further enrich ODEbench with more high-dimensional ODEs and provide a balanced dataset with over 156 samples. Each sample is accompanied by the variable descriptions and the domain name if available.
dx2/dt = (c8 + c9 * x3) - c10 * x2 x2: microg_dl
dx3/dt = (c4 + c5 * x1) / (1 + c6 * x2) - c7 * x3 x3: pg_ml
x2 = ( x1 AND NOT ( ( x3 OR x5 ) OR x4 ) ) x2: v_Emx2
x3 = ( ( x3 AND x5 ) AND NOT x2 ) x3: v_Fgf8
x4 = ( x5 AND NOT ( x2 OR x1 ) ) x4: v_Pax6
x5 = ( x3 AND NOT x2 ) x5: v_Sp8
$\sim$ x1: rapidly_activating_K_current_n_gate (dimensionless)
$\sim$ x2: Ca_i in component ionic_concentrations (micromolar)
$\sim$ x3: None
  1. (a)
    CDEs: We use an LLM to generate the skeleton of the continuous‐time dynamical system, replacing each unknown coefficient with a placeholder "c". An optimizer then fits these coefficients to the training time series. To avoid the high cost of repeatedly solving and differentiating through a full ODE solver, we adopt the finite–difference approximation strategy from ODEFormer d’Ascoli et al. (2023). Once the functional form $\hat{f}(\mathbf{x};\boldsymbol{\phi})$ is obtained, where $\phi$ is the fitted coefficients, we generate the final numerical solution using SciPy’s scipy.integrate.solve_ivp, and compare the trajectory to the ground truth:
    The distribution of sample dims is shown in Figure 7(a).
  2. (b)
  3. (c)
(b) Dimension distribution of Boolean Networks used in SymbolBench.
Appendix C LLMs for Benchmarking
We evaluate six representative LLMs, chosen for their diversity in size, training data, and specialization. For all models, we set the temperature=1, top-p=0.9, top-k=60, and the maximum tokens generated to be 4098 to explore diverse and creative symbolic structures.
  1. (a)
  2. (b)
  3. (c)
  4. (d)
  5. (e)
  6. (f)
Appendix D Baseline Implementation
  1. (a)
    CDEs: We follow ODEFormer’s baseline implementation and hyperparameter tuning d’Ascoli et al. (2023), using PySR with finite‐difference approximations for skeleton search, and default greedy top-$k$ generation for ODEFormer.
  2. (b)
    BNs: We reimplement LogicGep’s Boolean‐network inference using the same Geppy genetic‐programming framework Gao et al. (2022), but omit the continuous‐to‐binary discretization and MLP‐based constraint stages, since our training traces are already binary.
  3. (c)
    SCMs: All SCM baselines are based on Tigramite Runge et al. (2022), with the maximum time‐lag set to 1 for fair comparison.
Appendix E Full Results on the BN and SCM datasets
@@
Appendix F Error Analysis for Boolean Network Inference
Summarizing the performance on Boolean Network tasks, we categorize the errors from the following perspectives.
F.1 Limited Observations
Limited observation brings uncertainty in predicting the accurate form of the symbolic structure.
Example:
Complete Ground Truth Regulations:
x1 $=(\neg(x3\vee x5)\vee\neg(x5\vee x3))$ (2)
x2 $=(x1\wedge\neg((x3\vee x5)\vee x4))$ (3)
x3 $=((x3\wedge x5)\wedge\neg x2)$ (4)
Complete Predicted Regulations:
$x1$ $=(x2\vee\neg x3)\wedge(x4\vee\neg x5)$ (5)
$x2$ $=x2\wedge x3\wedge x5\wedge\neg x3$ (6)
$x3$ $=x3\wedge x5$ (7)
Analysis: This case demonstrates overfitting due to limited observations. The Boolean network has 5 variables with only 11 training transitions. The model achieves ID F1 = 0.571 but drops to OOD F1 = 0.087 (performance drop: 0.484). The predicted regulations show different logical structures that work for training data but fail to capture true Boolean network dynamics.
F.2 Aligning Symbolic Structure and Numerical Fitness
The numerical fitness in the OOD setting is largely determined by the symbolic structure accuracy. One difficulty during the iterative search process is that the proximity of symbolic structure is not correlated with the proximity of numerical fitness, which means numerically similar predictions can have a large difference in symbolic structure similarity.
Example:
Complete Ground Truth Structure:
x1 $=(((((((\neg x2\wedge\neg x4)\wedge x7)\vee(((\neg x2\wedge x4)\wedge\neg\ldots$ (8)
x2 $=(((((\neg x2\wedge\neg x4)\vee((\neg x2\wedge x4)\wedge x1))\vee((x2\wedge\ldots$ (9)
x3 $=((x2\wedge\neg x4)\vee((x2\wedge x4)\wedge x1))$ (10)
x4 $=((\neg x2\vee(x2\wedge\neg x3))\vee(((x2\wedge x3)\wedge x4)\wedge x5))$ (11)
x5 $=(((\neg x2\wedge x4)\vee((x2\wedge\neg x3)\wedge x4))\vee((((x2\wedge x\ldots$ (12)
x6 $=((((\neg x2\wedge x4)\wedge x5)\wedge\neg x6)\vee((((x2\wedge\neg x3)\wedge\ldots$ (13)
x7 $=((\neg x2\wedge\neg x4)\vee((\neg x2\wedge x4)\wedge\neg x5))$ (14)
Complete Predicted Structure:
$x1$ $=x7$ (15)
$x2$ $=x3\vee x4$ (16)
$x3$ $=x2\wedge x6$ (17)
$x4$ $=x6\vee x7$ (18)
$x5$ $=x3\vee x4$ (19)
$x6$ $=x5$ (20)
$x7$ $=x7$ (21)
Analysis: This case exemplifies structure-fitness misalignment where high NED = 17.43 indicates structural dissimilarity to ground truth, yet F1 performance is good (0.671). This demonstrates that functionally equivalent Boolean regulations can have very different symbolic representations. The model found an alternative solution that achieves similar numerical fitness despite having a completely different structural form. This highlights that tree edit distance may not capture functional equivalence in Boolean networks, where multiple valid regulatory patterns can produce equivalent behaviors.
F.3 Error in Context Alignment
While we use context to shrink the search space and hope to predict more accurate symbolic forms, a misalignment with the given context can produce the symbolic structure from the wrong domain.
Example:
Expected Boolean Network Structure:
x1 $=(\neg(x3\vee x5)\vee\neg(x5\vee x3))$ (22)
x2 $=(x1\wedge\neg((x3\vee x5)\vee x4))$ (23)
x3 $=((x3\wedge x5)\wedge\neg x2)$ (24)
x4 $=(x5\wedge\neg(x2\vee x1))$ (25)
x5 $=(x3\wedge\neg x2)$ (26)
Predicted Structure:
$x1$ $=(x2\vee\neg x3)\wedge(x4\vee\neg x5)$ (27)
$x2$ $=x2\wedge x3\wedge x5\wedge\neg x3$ (28)
$x3$ $=x3\wedge x5$ (29)
$x4$ $=x3\wedge x4$ (30)
$x5$ $=x4\wedge x5$ (31)
Biological Domain Context: This Boolean network models cortical area development during embryonic brain formation. The variables represent critical developmental genes with established biological functions: